OpenStudy (anonymous):

two negative integers have a sum of -7 and a product of 6 what are the integers

OpenStudy (anonymous):

By observation, the numbers are -1 and -6

OpenStudy (anonymous):

Thanks how would i solve this the width of a rectangle is 6lessthan its length and the area is 40sq inches what are the length and width

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Area of a rectangle = lb Given, b = l - 6 Substitute the value of b and you will get a quadratic equation. Solving the quadratic equation, you will get the values of l and b

OpenStudy (anonymous):

If you want a easier method, it can be observed that the factors of 40 are 10 x 4 and their difference is 6. So you can easily conclude that l = 10 and width = 4

OpenStudy (anonymous):

so if if the length of a rectangle plus its width is 25cm the area is 156 so to find the length and width im going 10 / 156by 4

OpenStudy (anonymous):

so it would be 4,39

OpenStudy (anonymous):

wt do u mean by "im going 10 / 156by"??

OpenStudy (anonymous):

am i going to divded 10by 156

OpenStudy (anonymous):

But 4 + 39 is not equal to 25 !!! :) The answer is 12 and 13

OpenStudy (anonymous):

I will explain you this one We have l+ b = 25 and lb = 156 We can write l= 25 - b and substitute the second equation and we get 12 and 13 as the answer